### v5

# The human-machine paradox:

Abstract
When deploying artificial skills, managers widely assume that combining them with the human factor is a safe harbor, mitigating the risks of full automation in high-complexity tasks. This paper formally challenges the economic validity of this widespread assumption, arguing that the true bottom-line economic utility of a human-machine skill policy is dangerously misunderstood and highly contingent on situational and design factors. To investigate this gap, we develop an in-silico framework based on Monte Carlo simulations grounded in empirical pragmatism to quantify the economic impact of human and machine skills in the execution of tasks presenting varying levels of complexity. Our results show that a human-machine strategy can yield the highest economic utility in complex scenarios, but only if genuine augmentation is achieved. In contrast, when failing to realize this synergy, the human-machine approach can perform worse than either the machine-exclusive or the human-exclusive policy, actively destroying value under the pressure of costs that are not compensated by sufficient performance gains. The takeaway for decision-makers is unambiguous: when the context is complex and critical, simply allocating human and machine skills to a task may be insufficient, and far from being a silver-bullet solution or a low-risk compromise. Rather, it is a critical opportunity to boost competitiveness that demands a strong organizational commitment to enabling augmentation. Also, our findings show that improving the cost-effectiveness of machine skills over time, while useful, does not replace the fundamental need to focus on achieving augmentation. By synthesizing task-based economic theory with the concept of generalization difficulty, our work offers a scalable and low-cost methodology for organizations to forecast the economic implications of skill policy decisions ex ante. This provides a foundation for scaling experiments and study more general policy implications as candidates for future empirical validation.
1 Introduction
Prominent research analyzing the impact of technology over skill demand shows that computer capital allocations tend to result in automating well-defined and rule-based routine tasks, complementing human skills in tasks requiring abstraction, problem-solving, or creativity (Autor et al., 2003), and supplementing human judgment (Autor, 2024), with experimental evidence in (Peng et al., 2023), (Noy and Zhang, 2023), and (Brynjolfsson et al., 2025), provided the ability to use artificial capabilities effectively (Agarwal et al., 2023) (Autor, 2024). Meanwhile, technological advancements in methods, combined with increasingly affordable computational capabilities, have allowed artificial skills to achieve remarkable performance in tasks where automation has historically struggled. Notable examples include image recognition (Krizhevsky et al., 2012), natural language processing (Vaswani et al., 2017), game play (Silver et al., 2017), skin cancer diagnostics (Esteva et al., 2017), recommendation systems (He et al., 2017) and protein-folding research (Jumper et al., 2021), just to mention some widely recognized examples. Moreover, language models deployed through deep neural networks (Brown et al., 2020) (Hoffmann et al., 2022) (Team et al., 2023) (Touvron et al., 2023) have contributed to amplify expectations regarding the present and future performance of artificial skills in the execution of a wide variety of cognitive tasks, making them extremely accessible to a broad audience through user-friendly interfaces. These examples emphasize the potential of technological progress to expand the set of tasks that can be automated (Acemoglu, 2025). However, when tasks are characterized by high levels of generalization difficulty Chollet (2019), machines can struggle to achieve high levels of performance, presumably due lack of the abstraction and analogy-making abilities typical of the human mind. Nevertheless, while human skills may outperform artificial skills when facing surprising task content, they may be less efficient and less scalable. Then, an alternative solution to the human-machine dichotomy is to combine these two skills to get the best of both, enhancing execution performance at a cost of the dual skill structure. From an organizational perspective, these conditions make the administrative behavior (Simon, 1947) required to determine optimal capital and labor allocations far from straightforward. We call these skill policy decisions, which are particularly interesting in case of cognitive tasks, because compartmentalizing them is harder than manual or routine ones. To analyze this problem, we combine a task-based framework from the theory of Economics with provisions from Computer Science and Statistics into a model suitable to generate synthetic data through Monte Carlo simulations. Then, we leverage synthetic data to analyze the economic effects of skill policy decisions, from task-execution performance to bottom-line utility, in relation with initial conditions. To this end, we assume that task performance is not a deterministic outcome, but an emergent property shaped by complex interactions between skills and tasks under varying levels of complexity. Our goal is to understand more about the factors that can influence the economic effects of skill policy decisions, by leveraging bottom-up simulation to quantify emerging insights that may be transparent to top-down analysis. In this context, the computer science literature has extensively documented the technical challenges of generalization and the performance limitations of models on out-of-distribution or complex tasks. Therefore, our contribution is not to simply rediscover this qualitative insight. We introduce a formal in-silico framework that translates these known technical performance characteristics into quantifiable economic utility measures. To our knowledge, no prior work provides a formal model for ex-ante evaluation of skill policies that explicitly bridges the technical concept of generalization difficulty with the economic concepts of error costs and bottom-line economic utility. Thus we genuinely believe that our proposed framework can be a novel and useful tool for decision-makers, expanding the discussion from skill performance alone to skill policy, and to how the skill policy of choice should be deployed to maximize economic utility.
2 Referential frameworks
Our work builds upon a task-based model of production (Farboodi and Veldkamp, 2021) characterizing the quality-adjusted output of a firm $i$ at time $t$ as $y_{i,t}=A_{i,t}k_{i,t}^{\alpha}$, where $k_{i,t}^{\alpha}$ represents the units of output of quality $A_{i,t}$ produced by leveraging $k_{i,t}$ units of capital, and $A_{i,t}=g\big((a_{i,t}-a_{i,t}^{opt})^{2}\big)$ is function of the squared distance between $a_{i,t}$ (the production technique of choice for firm $i$ at time $t$) and the optimal technique $a_{i,t}^{opt}=\theta_{t}+\epsilon_{a,i,t}$}, with $\theta_{t}$ and $\epsilon_{a,i,t}$ being the jointly-observed persistent and transitionary components of the optimal technique, and $g$ a strictly decreasing function. This formulation, by means of the introduction of the squared distance between the actual and the optimal production skill, makes the role of technological performance explicit in the determination of production output of automated tasks. Also, this framework complements the treatments illustrated in (Autor and Thompson, 2025), where a worker is able to execute a task only if the worker’s expertise level is equal or higher than the one required by the task (the hierarchical expertise assumption), and automation deploys below a certain threshold of task complexity (the hierarchical automation assumption), a condition also illustrated in (Acemoglu, 2025) as effect of the comparative advantage offered by labor with respect to capital in newer or complex tasks. Interestingly, this task-based model aligns with the idea of generalization difficulty as presented in (Chollet, 2019), i.e. the increased complexity introduced by additional levels of uncertainty presenting at execution time with respect to learning time. The compatibility of these seminal models constitutes the foundational ground for our work.
3 Model
By combining the provisions of the referential frameworks introduced in Section 2, we conceptualize performance as a stochastic property that emerges from the interaction between task content and skills, conditioned upon local and global generalization difficulty, extending the definition of production technique to include models operated by the human mind. In this context, we focus on cognitive tasks, leaving the opportunity to relax this assumption in further steps. Herein, we move by abstraction to defining the distance between the actual and optimal skills in terms of loss ($\mathcal{L}$), such that the quality $A_{i,t}$ can be expressed as $A_{i,t}=g\big(\mathcal{L}_{i,t}\big)$, where $\mathcal{L}_{i,t}=f(a_{i,t},a_{i,t}^{opt})$ and $f$ is any appropriate function measuring the absolute gap between the actual and optimal technological skills. By doing this, we open the possibility to consider alternative measures of distance, including divergence or dissimilarity (also, if necessary, multiple measures of loss may be combined). Then, when the range of potential values of $\mathcal{L}_{i,t}$ is known, we can obtain a normalized loss $\mathcal{L}^{\prime}_{i,t}\in[0,1]$ through the normalization process of choice (for example, $\mathcal{L}^{\prime}_{i,t}=[\mathcal{L}_{i,t}-\min(\mathcal{L}_{i,t})]/[\max(\mathcal{L}_{i,t})-\min(\mathcal{L}_{i,t})]$), and a normalized performance $\theta_{i,t}=(1-\mathcal{L}^{\prime}_{i,t})$. This allows to formally transition from the measure of technological distance to the measure of performance, with consequently easier operational deployment. Based on this initial framing, we implement a computational pipeline that, given initial conditions at time $t$, leverages a stochastic process to determine the normalized performance $\theta_{i,c,d,e}$ obtained by the firm $i$ deploying the skill policy $c$ in the execution of a task with difficulty $d$ at time $t+e$, where $e\in[0,E-1]$ is a task-execution epoch. To this end, without sacrificing generality, our design implements three skill policies and three levels of task difficulty. The skill policy options are the human-exclusive ($H$, relying on labor allocations only), the machine-exclusive ($M$, i.e. full automation), and the human-machine ($HM$, combining human and machine skills). The implementation of local generalization difficulty is probabilistic, with distributions $P(\theta)_{c,d,e}$ conditioned on $c$, $d$, and $e$. More specifically, the local generalization difficulty is inherently described by each individual probability distribution $P(\theta)_{c,d,e}$, while the global generalization difficulty is modeled by having different distributions for different levels of difficulty $d$. At each epoch, the performance obtained by the skill policies $H$ or $M$ (denoted as $\theta_{i,c=H,d,e}$ and $\theta_{i,c=M,d,e}$, and more synthetically as $\theta_{H}$ and $\theta_{M}$, respectively) is randomly drawn according to the appropriate probability distribution $P(\theta)_{c,d,e}$. When, instead, the skill policy adopted by the firm is $HM$, the performance computes as $\theta_{HM}=a(\theta_{H},\theta_{M},\gamma_{HM})$, where $a$ is an arbitrary function defining how the individual skills $H$ and $M$ interact, depending also on an augmenting factor $\gamma_{HM}$. Herein, the concept of augmentation evokes amplification (Ashby, 1956), collaboration (Licklider, 1960) and integration (Engelbart, 1962), characterizing a human-machine skill policy in ways that go from task decomposition and sub-task delegation to complex collaborative partnerships. Then, based on the performance $\theta_{i,c,d,e}$, the quality-adjusted production output computes as $y_{i,c,d,e}=g^{\prime}\big(\theta_{i,c,d,e}\big)$, where $g^{\prime}:\theta_{i,c,d,e}\in[0,1]\rightarrow y_{i,c,d,e}\in[0,1]$ is a monotone and non decreasing function (the combined evaluation of $A_{i,t}$ and ${kl}^{\alpha}_{i,t}$ through the function $g^{\prime}$ is a desired behavior of our model, to streamline the formal relationship between task-execution performance and quality-adjusted output). Next, the value of the production output is determined as $v_{i,c,d,e}=max(0,min(1,y_{i,c,d,e}*mc_{c}*\delta_{c,e}))$, where $mc_{c}\in[0,1],\forall c$ is the margin of contribution associated to each skill policy, with point values $mc_{H},mc_{M},mc_{HM}$, and $\delta_{c,e}$, which depends from both $c$ and $e$, represents the skill-specific evolution of the margin of contribution over epochs. To this end, $\delta_{c,e}$ is computed by linear interpolation as $\delta_{c,e}=interpolate(1,1+\delta_{c},e,E)$, according to Algorithm 3, where $\delta_{c}=\{\delta_{H},\delta_{HM},\delta_{M}\}$ are configuration parameters determining the increase or decrease of the margin of contribution over time, modeling the progress of the firm in walking the skill learning curve, or the firm’s ability to improve the cost effectiveness of skill development and operations, or the evolution of the market price for the task production under consideration ($\delta_{c}>0$ determines an increase, i.e. a unit of quality-adjusted production increases through epochs, while $\delta_{c}<0$ determines a decrease). Also, we introduce the element of error as a shock conditioned on performance, modeling the economic impact of a low performance not only in terms of reduced output, but also in terms of economic penalty triggered beyond a minimum performance threshold $t_{err}\in[0,1]$. This threshold can be interpreted as a quality control threshold, or a level of safety tolerance. In combination with $t_{err}$, we introduce the parameter $c_{err}\in[0,1]$, which defines the cost of error as a fraction of input in units ($c_{err}$). This represents a variety of phenomena, including the cost of the necessary corrections, the cost of repair for eventual damages, a financial penalty, a reputation damage, or the cost of a product recall. By combining these factors, when $(1-\theta_{i,c,d,e})\geq t_{err}$ we compute $err_{i,c,d,e}=(1-\theta_{i,c,d,e})*c_{err}$, or $err_{i,c,d,e}=0$ otherwise. Finally, the bottom-line economic utility is computed as $u_{i,c,d,e}=v_{i,c,d,e}-err_{i,c,d,e}$. The complete process instructed by the model is more formally describe by Algorithm 4, and schematically illustrated in Figure 1. As schematically illustrated in Figure 2, the result of each simulation $s$ is collected through the following sets of values: $\Theta^{s}=\{\theta_{i,c,d,e,k}^{s}\}$, $Y^{s}=\{y_{i,c,d,e,k}^{s}\}$, $V^{s}=\{v_{i,c,d,e,k}^{s}\}$, $ERR^{s}=\{err_{i,c,d,e,k}^{s}\}$, and $U^{s}=\{u_{i,c,d,e,k}^{s}\}$. For notational convenience, we generically refer to these sets as $\Lambda^{s}=\{\lambda_{i,c,d,e,k}^{s}\}$, with $\lambda\in\{\theta,y,v,err,u\}$ and $\Lambda\in\{\Theta,Y,V,ERR,U\}$. Additionally, we define the subsets of values characterized by the same skill policy $c_{0}$ as $\Lambda^{s}_{c_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0},\forall i,e,d,k\}\subseteq\Lambda^{s}$, and the subsets of values characterized by the same skill policy / task difficulty pair $(c_{0},d_{0})$ as $\Lambda^{s}_{c_{0},d_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0}\land d=d_{0},\forall i,e,k\}\subseteq\Lambda^{s}_{c_{0}}$. Next, for each set $\Lambda^{s}_{[c],[d]}$ (this compact notation stands for $\Lambda^{s}$, $\Lambda^{s}_{c}$ and $\Lambda^{s}_{c,d}$), we compute the arithmetic mean as $\mu(\Lambda^{s}_{[c],[d]})$, the standard deviation as $\sigma(\Lambda^{s}_{[c],[d]})$, the range as $\rho(\Lambda^{s}_{[c],[d]})=\max(\Lambda^{s}_{[c],[d]})-\min(\Lambda^{s}_{[c],[d]})$, the interquartile range (IQR) as $IQR(\Lambda^{s}_{[c],[d]})$, and the skewness as $SK(\Lambda^{s}_{[c],[d]})$. The descriptive statistics $\mu$, $\sigma$, $\rho$, $IQR$ and $SK$ are also generically denoted as $\omega$, crystallizing $\omega(\Lambda^{s}_{[c],[d]})$ as a general notational handles for the characterization and analysis of results.
Figure 1: Schematic representation of the computational flow implemented by the model
Algorithm 1 Function $randomchoice:(X,n,P(X))\rightarrow choice$
[…]
2:$n$
[…]
Return: choicechoice $\triangleright$ A vector of length $n$
Algorithm 2 Function $randomsample:(\alpha,\beta,n)\rightarrow sample$
[…]
3:$n$
[…]
Return: samplesample $\triangleright$ A vector of length $n$
Algorithm 3 Function $interpolate:(param\_start,param\_end,cur\_step,tot\_steps)\rightarrow cur\_param$
1:param_startparam\_start
2:param_endparam\_end
3:cur_stepcur\_step
4:tot_stepstot\_steps
5:if $total\_steps\leq 1$ then
6:  $cur\_param\leftarrow param\_start$
7:else
8:  $cur\_param\leftarrow param\_start+(param\_end-param\_start)*cur\_step/(tot\_steps-1)$
9:end if
Return: cur_paramcur\_param
Algorithm 4 Model
[…]
10:$K,N,E$
[…]
15:for $k=1$ to $K$ do
16:  for $e=0$ to E1E-1 do
17:    if $e=0$ then
18:     $c\leftarrow randomchoice(C,N,P_{c})$
19:    end if
20:    for $i=0$ to N1N-1 do
21:     $d\leftarrow randomchoice(D,1,P_{d})$
22:     $\alpha_{c,d,e}\leftarrow interpolate(\alpha_{c,d,e=0},\alpha_{c,d,e=E-1},e,E)$
23:     $\beta_{c,d,e}\leftarrow interpolate(\beta_{c,d,e=0},\beta_{c,d,e=E-1},e,E)$
24:     $\theta_{i,c=H,d,e,k}\leftarrow randomsample\big(\alpha_{c=H,d,e},\beta_{c=H,d,e},1)$
25:     $\theta_{i,c=M,d,e,k}\leftarrow randomsample\big(\alpha_{c=M,d,e},\beta_{c=M,d,e},1)$
26:     if $c=H$ then
27:       $\theta_{i,c,d,e,k}\leftarrow\theta_{i,c=H,d,e,k}$
28:     else if $c=M$ then
29:       $\theta_{i,c,d,e,k}\leftarrow\theta_{i,c=M,d,e,k}$
30:     else if $c=HM$ then
31:       $\theta_{i,c,d,e,k}\leftarrow a(\theta_{i,c=H,d,e,k},\theta_{i,c=M,d,e,k},\gamma_{HM})$
32:     end if
33:     $y_{i,c,d,e,k}\leftarrow g^{\prime}(\theta_{i,c,d,e,k})$
34:     $\forall c:\delta_{c,e}\leftarrow interpolate(1,1+\delta_{c},e,E)$
35:     $v_{i,c,d,e,k}\leftarrow y_{i,c,d,e,k}*mc_{c}*\delta_{c,e}$
36:     if $(1-\theta_{i,c,d,e,k})\geq t_{err}$ then
37:       $err_{i,c,d,e,k}\leftarrow(1-\theta_{i,c,d,e,k})*c_{err}$
38:     else
39:       $err_{i,c,d,e,k}\leftarrow 0$
40:     end if
41:     $u_{i,c,d,e,k}\leftarrow v_{i,c,d,e,k}-err_{i,c,d,e,k}$
42:    end for
43:  end for
44:end for
Returns: $\Lambda^{s}=\{\lambda_{i,c,d,e,k}^{s}\}$ with $\lambda\in\{\theta,y,v,err,u\}$, i.e. $\Theta^{s}=\{\theta_{i,c,d,e,k}^{s}\}$, $Y^{s}=\{y_{i,c,d,e,k}^{s}\}$, $V^{s}=\{v_{i,c,d,e,k}^{s}\}$, $ERR^{s}=\{err_{i,c,d,e,k}^{s}\}$, $U^{s}=\{u_{i,c,d,e,k}^{s}\}$ with $s$ denoting the identifying index of the simulation of interest.
Figure 2: Schematic representation of an experiment combining multiple simulations with different input parameters, and the sub-setting/processing of the resulting data
4 Design of experiment
In our Monte Carlo experimental setting, we focus on a single firm ($N=1$) walking through ten epochs ($E=10$). With regards to the number of replications, which remarkably impacts the computational effort required by simulations, we initially set $K=1000$, checking the margin of error at 95% confidence level before evaluating an eventual increase of $K$. Then, for the function $a$, we implement three options:
$\theta_{HM}=\begin{cases}(\theta_{H},\theta_{M})/2&\text{if }a=mean,\\ \min(1,mean(\theta_{H},\theta_{M})*\gamma_{HM}))&\text{if }a=collaborate,\\ \min(1,max(\theta_{H},\theta_{M})*\gamma_{HM}))&\text{if }a=superpower,\end{cases}$ (1)
where $\theta_{HM}$ is a notational simplification for $\theta_{i,c=HM,d,e},\forall(i,d,e)$. Herein, meanmean represents the case of simply averaging the abilities of human and machine skills with no augmenting effects, while collaboratecollaborate and superpowersuperpower represent moderate and strong augmentation, respectively. For the function $g^{\prime}$, we deploy two options:
$y=\begin{cases}1/(1+e^{-k(2\theta-1)})&\text{if }g^{\prime}=logistic,\\ [k-\ln(1-\theta)-\ln(\theta)]/2k&\text{if }g^{\prime}=inverse\_logistic,\end{cases}$ (2)
where $y$ is a notational simplification for $y_{i,c,d,e}$ and $\theta$ is a notational simplification for $\theta_{i,c,d,e}$. These options reflect two contrasting dynamics, offering performance-to-output maps. Then, we set $\gamma_{HM}\in[1,2]$, thus limiting the augmenting factor to increase only, in the range [0%,100%]. Also, we impose $mc_{H}\in[0.4,0.6]$ and $mc_{M}\in[0.7,0.9]$, thus putting us in the case of automation bringing higher margin of contribution than human labor, and $mc_{HM}=1-(1-mc_{H})-(1-mc_{M})$ to represent the idea that both labor and capital allocations are required when deploying human-machine collaboration. Finally, we set $\delta_{H}=0$ and $\delta_{M}=0$ as static default, representing the case of human and machine skills being already on their plateau of efficiency. According to these initial constraints, we formulate our experimental design for numeric variables by combining the value selection optimized through the Latin Hypercube Sampling method (LHS) (McKay et al., 2000) and the Maximin method (Johnson et al., 1990) -see Figure 3 for an overview of the resulting pairwise value selection. With regards to the distributions of skill policies and task difficulties, we set $P_{c}\sim U\{P_{H},P_{HM},P_{M}\}$ and $P_{d}\sim U\{P_{Low},P_{Med},P_{High}\}$, i.e. uniformly distributed, and we opt for $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$ for the probability distributions of performance. More in depth, in a first experiment we set the parameter pairs $(\alpha,\beta)_{c,d,e=0}$ and $(\alpha,\beta)_{c,d,e=E-1}$ configuring improvement for human skills across all task difficulties, with static machine skills, as illustrated in Figure 4(a). In a second experiment, which is meant to stress test the simulation results under the effect of improved machine skills, we set the parameter pairs $(\alpha,\beta)_{c,d,e=0}$ and $(\alpha,\beta)_{c,d,e=E-1}$ as illustrated in Figure 4(b). All the experiments are executed on computing instances deploying Python 3.11, generating a total of 57.6 million task executions and 288 million data logs for $\lambda_{i,c,d,e,k}$ with $\lambda\in\{\theta,y,v,err,u\}$.
(a) LHS method
(b) Maximin method
Figure 3: Pair plot of numeric variables selected for experimental design instructed by optimization methods
(a) Exp. 1
(b) Exp. 2
Figure 4: Distributions of probabilities $P(\theta)_{c,d,e}$ for all skill policies and difficulty levels at epoch $e=0$ and e=E1e=E-1, representing human skills to evolve through epochs, while machine skills are static
5 Results
Contrary to the conventional wisdom that a human-in-the-loop policy is a safe risk-mitigating strategy, our simulations reveal it prone to contextual economic volatility. As shown in Figure 5, in tasks requiring high generalization capabilities, the human-machine skill policy over performs both the human-exclusive and the machine-exclusive strategies in terms of bottom-line utility, but only when achieving augmenting effects (central tendency of mean utility up to 0.22 for the human-machine team vs 0.08 and 0.03 for the human and machine alone, respectively). Instead, when failing to produce augmenting effects, the human-machine strategy often results the least economically effective, sometimes entering the dangerous zone of negative utility, destroying value instead of building it. Interestingly, the advantage of the human-machine skill policy achieving augmenting effects in high-difficulty tasks is confirmed also when machine skills surpass the performance of human skills across all task difficulties if the context is characterized by high cost of errors (experiment 2, with $c_{err}\geq 0.80$ and $t_{err}\leq 0.10$). While denoting the emergent critical finding regarding the all-or-nothing nature of the human-machine skill policy, our results successfully replicate two relevant established stylized facts from the existing literature. First, the simulations show that automation tends to be the dominant economic strategy when task difficulty is low-to-medium, underscoring its superiority for repetitive or compartmentalized tasks. Second, the framework confirms that the human-exclusive skill policy can be highly competitive, and often superior, when surprises at inference time are the norm and generalization difficulty is high.
(a) Performance ($\Lambda=\Theta$)
(b) Utility ($\Lambda=U$)
Figure 5: Mean of the distributions $\mu(\Lambda^{s}_{c,d})$ with $\Lambda\in\{\Theta,U\}$, with breakdown by experiment and by human-machine interaction dynamics (function $a$)
Moving from overview to a more fine-grained detail, we consolidate three characterizations of our results, with distinction between Experiment 1 and 2, including the interesting close-up on the data of Experiment 2 with focus on the conditions characterized by high cost of error. First, in Table 1 we present a synthetic view over the central tendency for the distributions $\omega(\Lambda^{s}_{[c],[d]})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ under the assumption of uniformly-distributed probability of occurrence for configuration setups, with incremental breakdown by skill policy and task difficulty. Then, in Table 2, we extend the statistical description of these distributions by reporting their standard deviation, range, and median value; additionally, we provide the breakdown by human-machine interaction dynamics (function $a$) for the human-machine skill policy. Next, we visualize the distributions $\omega(\Lambda^{s}_{[c],[d]})$ with $\omega\in\{\mu,\sigma\}$ and $\Lambda\in\{\Theta,Y,V,ERR,U\}$ based on overlapping violin and box plots, (see Figures 6 and 7).
Table 1: Central tendency (arithmetic mean) of the distributions $\omega(\Lambda^{s}_{[c],[d]})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$
Table 2: Statistical description (mean, standard deviation, range, and median) of the distributions $\omega(\Lambda_{[c],[d]}^{s})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$
(a) Exp. 1
(b) Exp. 2
(c) Exp. 2*
Figure 6: Visualization based on overlapping violin and box plots of the distributions $\mu(\Lambda_{[c],[d]}^{s})$ with $\Lambda\in\{\Theta,Y,V,ERR,U\}$ by experiment
(a) Exp. 1
(b) Exp. 2
(c) Exp. 2*
Figure 7: Visualization based on overlapping violin and box plots of the distributions $\sigma(\Lambda_{[c],[d]}^{s})$ with $\Lambda\in\{\Theta,Y,V,ERR,U\}$ by experiment
6 Discussion
The visualization of the distributions $\mu(\Theta^{s}_{[c],[d]})$ and $\mu(U^{s}_{[c],[d]})$ resulting from both our experiments (see Figure 8) complements the overall effect revealed by the analysis of central tendencies. When the machine skills achieve remarkable performance across all task difficulties (Exp. 2), a straightforward effect is an overall dominance of the machine-exclusive skill policy in terms of utility, with consequential benefits also for the human-machine skill policy -see Figure 8(e). However, as detailed in Figure 8(f), which focuses the lens on situations characterized by potentially high cost of error, when task difficulty is high the human-machine skill policy achieving augmentation still prevails in terms of utility over the machine-exclusive policy. This reinforces the idea that the understanding of the context is crucial for optimal skill policy decisions. Moreover, when narrowing the analysis to contrasting scenarios of cost effectiveness, we observe that an increase over epochs of the margin of contribution associated to the human-machine skill policy mitigates the erosion of value, while not safely resolving it (see Figure 9, with the human-machine skill policy failing to achieve augmentation closing just at break-even). This highlights the importance of the augmenting effect towards the bottom-line economic effectiveness of the human-machine skill policy, even when machine skills do a great job in high-difficulty tasks. With regards to dispersion, the mean of the distributions $\sigma(U^{s}_{[c],[d]})$ and $\rho(U^{s}_{[c],[d]})$ presented in Figure 10 highlights a lower variability for the values in $U^{s}_{[c],[d]}$ over the simulation space. In contrast, in Figure 11 we observe that the standard deviation and the range of the distribution $\mu(U^{s}_{[c],[d]})$ for the human-machine skill policy tends to be slightly higher than than the one realized by the human or machine skills alone. Herein, on both the observations, we note that augmenting effects seem not to have a visible impact. Then, our interpretation is that the hybrid skill strategy contributes more to reducing the variability of utility over the spectrum of scenarios under consideration, while slightly but less evidently increasing the variability of $\mu(U^{s}_{[c],[d]})$. This may explain why the combination of human and machine skills can be perceived as a risk-mitigating strategy (the reduction of variability naturally suggests lower risks); however, the variability alone is not sufficient to avoid the risk of being near or below utility break even. To allow the reader a more precise understanding of the margins of error of $\mu(\Theta^{s}_{[c],[d]})$ and $\mu(U^{s}_{[c],[d]})$ obtained with $K=1000$ as estimates of the true mean ($K\rightarrow\infty$), for each simulation $s$ we compute the estimated MOE at 95% confidence level for $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$ and $\mu(\Lambda^{s}_{c,d})$ with $\Lambda\in\{\Theta,U\}$. Formally, $\forall s$, we compute $\text{MOE}_{95}\big(\mu(\Lambda^{s})\big)=1.96*\sigma(\Lambda^{s})/\sqrt{K^{s}}$, $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c})\big)=1.96*\sigma(\Lambda^{s}_{c})\sqrt{K^{s}_{c}}$, and $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c,d})\big)=1.96*\sigma(\Lambda^{s}_{c,d})\sqrt{K^{s}_{c,d}}$ with $K^{s}=K*E=10000$, and $K^{s}_{c}$ and $K^{s}_{c,d}$ globally and prudentially set to $3000<K*E/|C|\sim 3333$ and $1000<K*E/(|C|*|D|)\sim 1111$, respectively. We acknowledge that, within each simulation $s$, the sets $\Lambda^{s}$, $\Lambda^{s}_{c}$ and $\Lambda^{s}_{c,d}$ pool values that are related to different epochs $e$, thus potentially in relationship with different Beta distributions for skill policies $H$ and $HM$ (this is due to the deterministic interpolation of the parameters $\alpha$ and $\beta$ over epochs for the skill $H$), meaning that the data is heterogeneous and consists of independent samples of smaller size. However, the margins of error computed on the pooled standard deviation represent a conservative upper bound for the granular error that should be selectively computed for each epoch, because the pooled standard deviation combines the actual sampling uncertainty with the additional variability introduced by the parameter interpolation. The distributions of $\text{MOE}_{95}\big(\mu(\Lambda^{s})\big)$, $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c})\big)$ and $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c,d})\big)$ are described in Figure 13 and in Table 13, while Figure 13 presents the the means $\mu(\Lambda^{s}_{[c],[d]})$ against the related $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{[c],[d]})\big)$.
(a) $\mu(\Theta^{s}_{[c],[d]})$, Exp. 1
(b) $\mu(\Theta^{s}_{[c],[d]})$, Exp. 2
(c) $\mu(\Theta^{s}_{[c],[d]})$, Exp. 2*
(d) $\mu(U^{s}_{[c],[d]})$, Exp. 1
(e) $\mu(U^{s}_{[c],[d]})$, Exp. 2
(f) $\mu(U^{s}_{[c],[d]})$, Exp. 2*
Figure 8: Distributions of the values in the data sets $\mu(\Lambda^{s}_{[c],[d]})$ with $\Lambda\in\{\Theta,U\}$ with breakdown by human-machine interaction dynamics (function $a$), influencing augmenting effects
Figure 9: Experiment 1, mean of the distribution $\mu(U^{s}_{c,d})$, with breakdown by human-machine interaction dynamics (function $a$). The bar plots titled Exp. 1#1\# and Exp. 1£ show the mean of the distribution $\mu(U^{s}_{c,d})$ computed on subsets of $U^{s}_{c,d}$ data with $\delta_{HM}\leq 0.2$ and $\delta_{HM}\geq 0.8$, respectively. In short, Exp. 1#1\# focuses on low growth of margin of contribution over epochs for human-machine skills, while Exp. 1£ focuses on high growth.
(a) Mean of the distributions $\sigma(U^{s}_{c,d})$
(b) Mean of the distributions $\rho(U^{s}_{c,d})$
Figure 10: Mean of the distributions $\omega(U^{s}_{c,d})$ with $\omega\in\{\sigma,\rho\}$, with breakdown by experiment and by human-machine interaction dynamics (function $a$)
(a) Standard deviation of the distributions $\mu(U^{s}_{c,d})$
(b) Range of the distributions $\mu(U^{s}_{c,d})$
Figure 11: Standard deviation and range of the distributions $\mu(U^{s}_{c,d})$, with breakdown by experiment and by human-machine interaction dynamics (function $a$)
(a) $\Lambda^{s}$
(b) $\Lambda^{s}_{c}$
(c) $\Lambda^{s}_{c,d}$
Figure 12: Histograms of the distributions of the margins of error for $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$ and $\mu(\Lambda^{s}_{c,d})$ with $\Lambda\in\{\Theta,U\}$
(d) $\Lambda^{s}$
(e) $\Lambda^{s}_{c}$
(f) $\Lambda^{s}_{c,d}$
Figure 13: Scatter plots of $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{[c],[d]})\big)$ versus $\mu(\Lambda^{s}_{[c],[d]})$ with $\Lambda\in\{\Theta,U\}$
Table 3: Descriptive statistics for $\text{MOE}_{95}(\Lambda^{s}_{[c],[d]})$ with $\Lambda\in\{\Theta,U\}$
$\text{MOE}_{95}\big(\mu(\Theta^{s})\big)$ $\text{MOE}_{95}\big(\mu(U^{s})\big)$ $\text{MOE}_{95}\big(\mu(\Theta^{s}_{c})\big)$ $\text{MOE}_{95}\big(\mu(U^{s}_{c})\big)$ $\text{MOE}_{95}\big(\mu(\Theta^{s}_{c,d})\big)$ $\text{MOE}_{95}\big(\mu(U^{s}_{c,d})\big)$
mean 0.004 0.005 0.007 0.007 0.009 0.009
std 0.001 0.002 0.002 0.004 0.003 0.006
IQR 0.001 0.002 0.002 0.006 0.005 0.008
max 0.005 0.011 0.010 0.024 0.014 0.037
Q3 0.005 0.006 0.007 0.009 0.011 0.013
Q2 0.004 0.005 0.007 0.006 0.009 0.008
Q1 0.004 0.004 0.006 0.004 0.007 0.004
min 0.003 0.002 0.001 0.000 0.000 0.000
Finally, to better understand the importance of each framework’s variable on bottom-line utility, we leverage introspection analysis on meta-models learned on our synthetic data by means of i) Permutation Importance analysis and ii) SHapley Additive exPlanations method (SHAP). Permutation Importance analysis (Breiman, 2001) is a computational method measuring the importance of a feature by determining the increase of the prediction error after the values of that feature are randomly permuted (a larger increase in error denotes a more important feature). In the context of our work, we apply Permutation Importance analysis on each of the models trained in cross validation, finally computing mean and standard deviation of the results; also, we aggregate categorical values encoded in one-hot variables in a unique feature, improving interpretability. In contrast, the SHAP method (Štrumbelj and Kononenko, 2014) (Datta et al., 2016) (Lundberg and Lee, 2017), named by the acronym of SHapley Additive exPlanations, computes the marginal contribution provided by each feature to the output for any given example, providing local interpretability. Then, global model explanations are obtained by aggregating Shapley values across all examples in the data set, suggesting which features are most influential overall, and how they impact the model’s predictions. As the features $c$ and $a$ have a strong correlation by construction, to avoid compromising the response of permutation analysis and SHAP (Strobl et al., 2008) (Lundberg and Lee, 2017), we combine these features in a unique new variable called c_ac\_a, which jointly represents skill policy and augmenting effects. According to these provisions, we learn two models: $\hat{f}:(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})\rightarrow\mu(U_{c,d})$ and $\hat{f}:(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})\rightarrow\mu(U_{c})$. The absence of correlation between each models’ variable is verified by computing their adjusted Generalized Variance Inflation Factor (Adj.GVIF) (Fox and Monette, 1992), which results consistently in the range $[1.00,1.01]$ (no correlation). With regards to the modeling strategy, we implement Random Forest regressors (RFr) and Gradient Boosting regressors (GBr), assessing their performance through 10-Fold Cross Validation (metrics: $MSE$ and $R^{2}$, see Table 4; the purpose of measuring the predictive performance is only to make sure that models are effectively compressing data and learning the information encoded in it, an essential pre-requisite for the interpretation of variable importance). The results illustrated in Figure 14 and in Figure 15 suggest that the skill/augmentation policy c_ac\_a is the most important variable in terms of effect on bottom-line utility. Naturally, towards $\mu(U^{s}_{[c],[d]})$, also task difficulty $d$ is a remarkably important variable, increasingly when the distributions $P(\theta)_{c,d,e}$ differ upon changes in $d$ (the setup of the experiments is discriminant for the emergence of this effect). Then, a group of variables emerges for a second-level importance, including the performance threshold triggering the error shocks ($t_{err}$), the cost of errors ($c_{err}$), the baseline margin of contribution associated to the human-machine skill policy ($mc_{HM}$) and the performance-to-output function $g^{\prime}$. In contrast, the augmenting factor ($\gamma_{HM}$) shows to play a minor role, which suggests that the augmenting gain may be less important than the dynamics of skill collaboration and interplay.
Table 4: Combined results of the 10-Fold Cross Validation (MSE, $R^{2}$) for Exp. 1 and Exp. 2
Exp. 1 Exp. 2
RFr GBr RFr GBr
Meta-model MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$
$\mu(U_{c,d})=\hat{f}(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$ 0.000 0.994 0.004 0.932 0.000 0.996 0.003 0.945
$\mu(U_{c})=\hat{f}(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$ 0.000 0.991 0.002 0.925 0.000 0.996 0.002 0.957
(a) Exp. 1, RFr
(b) Exp. 1, GBr
(c) Exp. 2, RFr
(d) Exp. 2, GBr
(e) Exp. 1, RFr
(f) Exp. 1, GBr
(g) Exp. 2, RFr
(h) Exp. 2, GBr
Figure 14: Aggregated Permutation Importance analysis. The first row shows the mean decrease of $R^{2}$ for the model $\mu(U_{c,d})=\hat{f}(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$, while the second row relates to the model $\mu(U_{c})=\hat{f}(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$.
(a) Exp. 1, RFr
(b) Exp. 1, GBr
(c) Exp. 2, RFr
(d) Exp. 2, GBr
(e) Exp. 1, RFr
(f) Exp. 1, GBr
(g) Exp. 2, RFr
(h) Exp. 2, GBr
Figure 15: SHAP summary (beeswarm) plot illustrating variables’ impact on the output of models. The first row shows the mean decrease of $R^{2}$ for the model $\mu(U_{c,d})=\hat{f}(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$, while the second row relates to the model $\mu(U_{c})=\hat{f}(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$. Variables are ranked by global importance. Each point represents the SHAP value for a specific instance, with the color indicating the feature’s value. Categorical values (extremes): for task difficulty $d$, blue dots correspond to Low and red dots to High; for skill/augmentation policy c_ac\_a, blue dots correspond to H individual and red dots to HM superpower, with intermediate color scale passing through HMHM mean, $M$ individual (grey), and HMHM collaborate; for the function $g^{\prime}$, the blue dots represent logistic, while red dots represent inverse_logistic.
7 Limitations
Combining Monte Carlo simulations with machine learning meta-models can provide valuable insights, but limitations must be acknowledged to contextualize findings appropriately. The primary limitation of this study is that its conclusions are derived from synthetically generated data. Essentially, the simulation creates an idealized in silico environment whose potential gaps with the real-world should always be considered with the highest regard, as complex or unmodeled interactions could fundamentally alter the outcomes (Law et al., 2007). As a consequence, the validity of our findings is contingent upon the fidelity of our model, which is governed by two sets of assumptions: design hypotheses, related to the theoretical mechanics and interactions encoded in the model, and setup hypotheses, concerning the specific characterization of the scenarios of interest through parameters and probability distributions chosen for the stochastic elements. Then, before verification and validation, results and insights should be intended as specific to the explored parameter space. Design hypotheses include the specific choice in terms of distributions (for example, the choice of $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$) and the logics of computation of economic measures (for example, the cost of errors). Another arbitrary design choice pertains the specific integration of stochastic effects, which in our work are limited to the task-execution performance, leaving the other measures of economic impact to deterministic computation. This can potentially inject a re-scaling effect down through the funnel of economic value, thus making the dynamics of bottom-line measures (for example, utility) more a direct reflection of performance rather than an emergent behavior. With regards to setup hypotheses, while our experimental setup parameters are designed for real-world plausibility, the model’s external validity ultimately requires empirical testing. With regards to introspection, we underline that Permutation Importance and SHAP are not meant to reveal a causal relationship between features and outcome, but importance with respect to the predictions computed by the model under consideration. Moreover, we recognize that our definition of utility is focused on economics, thus representing a potentially incomplete picture when seen from another perspective (for example, from the point of view of social sciences). A final consideration pertains the choice of the number of replications $K$ for each simulation, which trades off the processing effort required by the experiment for the estimated margin of error (MOE) related to $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$, and $\mu(\Lambda^{s}_{c,d})$ computed over the replication space. While a number of replications between 5000 and 10000 should be recommended (Mundform et al., 2011), one of our goals is to minimize the computational effort to streamline the time and cost of simulations given an acceptable margin of error. Therefore, the considerations based on the analysis of $\mu(\Lambda^{s}_{[c],[d]})$ are conditioned upon the acceptance of the levels of MOE estimated in Section 6, Figures 13 and 13, and in Table 13. For the purpose of our study, we acknowledge the current margin of error as sufficient. Should a lower margin of error be required by the reader, the number of replications $K$ would need to be increased. To this end, our code is available with MIT license upon request.
8 Conclusions
The common wisdom in the deployment of artificial skills is built on a comforting assumption: combining humans and machines is a defensive safeguard against worst-case scenarios. This policy is believed to mitigate the risks of full automation while capturing some of its benefits by means of a moderate and cautious compromise. This paper formally challenges this assumption, arguing that it may be not only incorrect, but also economically dangerous, thus paradoxical. Our findings reveal that, in high-difficulty tasks, the human-machine skill policy is a high-risk, high-reward strategy. This is the primary contribution of our work. We show that in tasks characterized by high generalization difficulty, when genuine augmentation is not achieved, the human-machine approach frequently becomes the worst of all possible options, and an economic pitfall that actively destroys value by incurring the costs of both human and machine components without a synergistic benefit. In many common scenarios, a manager would be economically better off choosing either a human-only or a machine-only policy. However, the risk is mirrored by an equally high potential reward. When surprise is the norm, if genuine augmentation is achieved, the combination of human and artificial skills can unlock the highest economic utility, outperforming either humans or machines alone. This finding fundamentally highlights the two-folded nature of the managerial challenge of the AI era. The first strategic decision pertains whether or not to blend human and artificial skills. On this subject, our study suggests that the solution to this problem is highly contextual, depending on the skills available, on their performance, and on the difficulty of the task. The second organizational challenge is how to combine human and machine skills towards augmentation. Herein, the default human-in-the-loop paradigm may be insufficient, requiring a focused design effort to develop a collaborative partnership.
The in-silico framework we propose is a direct-to-purpose tool for ex-ante validation, allowing leaders to test the economic viability of their skill policies before costly implementation. By quantifying the economic utility of different skill policies, our experiments provide the foundation for a structured managerial approach. This framework moves beyond static task allocation and instead proposes a dynamic process for effective skill policy evaluation articulated in four stages:
  1. 1.
    Economic problem decomposition (understand the puzzle). Before a policy can be evaluated, the managerial problem must be formally decomposed. This requires to i) rigorously define tasks, and classifying them by generalization difficulty, ii) model task-execution performance in terms of distribution of probability for all available skills relevant to the task, and iii) map the economic utility of outcomes, including a clear quantification of the cost of errors;
  2. 2.
    Focus on synergistic augmentation (design 1+1=3). Our findings suggest that the human-machine policy is a high-variance strategy whose success is contingent on achieving genuine augmentation. This implies that managers must shift their focus from deciding allocations to designing collaborative workflows that actively engage human and machine skills in a process that can allow a better handling of generalization difficulty;
  3. 3.
    Policy simulation (search at scale). Given the complex, emergent nature of skill-task interactions, a priori analytical solutions are often intractable. The methodology presented in this paper offers means of searching at scale. By calibrating the framework with the parameters and methods determined at stage 1 and 2, decision-makers can conduct in-silico experiments to explore the economic outcome of competing skill policies under various conditions, identifying the option that yields the maximum expected utility, while mitigating downside risks;
  4. 4.
    Dynamic policy refinement (act and adapt). The method we propose is not a static, one-time tool but a mechanism for adaptive management. Data gathered from real-world executions of tasks should be used to continuously re-calibrate the model. This creates a dynamic feedback loop, enabling organizations to refine and adapt their skill policies in response to evolving technological capabilities and changing economic conditions.
Ultimately, our framework provides a quantitative method for organizations to move from general heuristics to data-driven policy decisions, allowing for a nuanced test-before-invest evaluation of the trade-offs between automation, human expertise, and human-machine collaboration. Naturally, this study has limitations: while our framework is a simulation grounded in empirical pragmatism, it is not a direct field study. Therefore, empirical testing is the great next step. Additionally, future work could extend this model to include dynamic learning and the impact of different organizational structures. This paper serves as a formal economic challenge to the myth of human-machine integration as a safe harbor. It is a call to move beyond wishful thinking and toward a critical, evidence-based approach for deciding when (and when not) to combine human and artificial skills, improving the organization’s competitiveness and the economic effectiveness of AI initiatives through highly-effective skill policy decisions.
9 Future developments
Empirical validation would radically strengthen this work, and should be the next-step priority. Additionally, modeling the unitary margins of contribution and the cost of error as probability distributions to reflect the variability of cost structures may be considered. Similarly, exploring other setups for the parameters characterizing the probability distributions of performance may be a promising option, together with considering a variety of classes of distributions. Moreover, the evolution of skills through linear interpolation may be refined to consider the nuances of skill learning and adaptation. Also, modeling the evolution of skills as a non-linear function of cumulative experience, or data generated, is an interesting perspective, in line with the provisions illustrated in (Farboodi and Veldkamp, 2021), where data is a defined as a by-product of production potentially enabling and feeding a feedback loop. Finally, future designs may specifically model dynamic training for machine skills, to account for the differences distinguishing the training of statistical models from heuristics or, more in general, of the models operated by the human mind.
10 Acknowledgments
The author declares no conflict of interest and no competing interests at the time this article is published.