### v6

# The human-machine paradox: why a dual skill strategy may not be the safest harbor when surprise is the norm, and why focusing on augmentation is key to resolving it

Abstract
When deploying artificial skills, decision-makers often assume that layering human oversight is a safe harbor that mitigates the risks of full automation in high-complexity tasks. This paper formally challenges the economic validity of this widespread assumption, arguing that the true bottom-line economic utility of a human-machine skill policy is dangerously misunderstood and highly contingent on situational and design factors. To investigate this gap, we develop an in-silico exploratory framework for policy analysis based on Monte Carlo simulations to quantify the economic impact of skill policies in the execution of tasks presenting varying levels of complexity across diverse setups. Our results show that in complex scenarios, a human-machine strategy can yield the highest economic utility, but only if genuine augmentation is achieved. In contrast, when failing to realize this synergy, the human-machine approach can perform worse than either the machine-exclusive or the human-exclusive policy, actively destroying value under the pressure of costs that are not sufficiently compensated by performance gains. This finding points to a key implication for decision-makers: when the context is complex and critical, simply allocating human and machine skills to a task may be insufficient, and far from being a silver-bullet solution or a low-risk compromise. Rather, it is a critical opportunity to boost competitiveness that demands a strong organizational commitment to enabling augmentation. Also, our findings show that improving the cost-effectiveness of machine skills over time, while useful, does not replace the fundamental need to focus on achieving augmentation when surprise is the norm, even when machines become more effective than humans in handling uncertainty. By synthesizing task-based economic theory with the concept of generalization difficulty, our work offers a scalable and low-cost methodology for organizations to forecast the economic implications of skill policy decisions ex ante. This provides a foundation for scaling experiments and study more general policy implications as candidates for future empirical validation.
1 Introduction
Research shows that technology can successfully automate routine tasks and complement human skills in more complex work (autor2003skill), while increasingly supplementing human judgment in cognitive tasks (autor2024applying), with supporting experimental evidence in (peng2023impact), (noy2023experimental), and (brynjolfsson2025generative). In this context, organizations of all kinds have a clear opportunity: leverage machine skills to enhance their capabilities and increase competitiveness. Nevertheless, true success hinges not simply in the adoption of new technologies, but on their effective use (agarwal2023combining) (autor2024applying). In turn, this depends not only on the performance achieved by individual skills, whether artificial or human, but also on how they are combined, and on how their use converts into net economic value. However, human and machine skills have distinct characteristics and different performance outcomes across various conditions, as well as different cost and margin structures. This disparity makes the the administrative behavior (Simon1947AdministrativeBehavior) required to determine optimal capital and labor allocations very complex, and risky. We call this the skill policy decision problem, whose solution under uncertainty is far from straightforward.
To understand its underlying nuances, one must first consider the modern evolving landscape of machine capabilities versus their traditional limitations. Recent advancements in methods, combined with increasingly affordable computational capabilities, have allowed machine skills to achieve remarkable performance in tasks where automation has historically struggled. Notable examples include image recognition (krizhevsky2012imagenet), natural language processing (vaswani2017attention), game play (silver2017mastering), skin cancer diagnostics (esteva2017dermatologist), recommendation systems (he2017neural) and protein-folding research (jumper2021highly), just to mention some widely recognized examples. Moreover, models of language deployed through deep neural networks (brown2020language) (hoffmann2022training) (team2023gemini) (touvron2023llama) have contributed to amplify expectations regarding the present and future performance of machine skills in the execution of a wide variety of cognitive tasks. Nevertheless, while machines have proved to excel in learning models on complex and vast data, “crush[ing] the barrier of meaning” (rota1985barrier) in tasks presenting high generalization difficulty (chollet2019measure) can still be challenging for automation. Achieving generalization under uncertainty is a non-monolithic ability involving human-like abstraction and analogy-making (mitchell1993analogy), which can be extremely hard for skill programs to replicate algorithmically. This explains why, as technology improves, the set of tasks that can be automated tends to expand (acemoglu2025simple), but not without limits. Beyond them, in highly uncertain or surprising situations, human skills are often preferred, paying a price in terms of lower efficiency and scalability. In this condition, and especially in the context of cognitive tasks (which are harder to compartmentalize), a potential and rather conventional solution is to avoid the skill policy dilemma altogether and combine human and machine skills to get the best of both. But is this always the safe harbor it seems?
To tackle this question, we build a generative model to analyze the economic effects of skill policy decisions across varying levels of generalization difficulty, investigating the sufficiency of plausible micro-foundational economic assumptions in relation to the emergence of the human-machine paradox. To accomplish this, we combine a task-based framework from the theory of Economics with provisions from Computer Science and Statistics into a model suitable to generate synthetic data through Monte Carlo simulations. This in-silico framework, designed for theoretical explanation and policy exploration (epstein2008model) rather than for empirical prediction, employs bottom-up simulation to quantify emergent insights. The framework operates on the assumption that performance is not a deterministic outcome, but an emergent property shaped by complex interactions between skills and tasks under varying levels of complexity. While the technical challenges of generalization are well-documented in computer science, this work’s primary contribution is not to rediscover that insight. Instead, it introduces a formal method to translate the generalization problem into quantifiable measures of economic utility. To our knowledge, no prior model offers an ex-ante evaluation of skill policies by explicitly bridging the technical concept of generalization difficulty with the economic concepts of error costs and utility. Therefore, this framework is proposed as a novel tool for decision-makers, expanding the discussion from skill performance alone to skill policy decisions, and to how the skill policy of choice needs to be deployed to maximize economic utility.
2 Model
Our work builds upon a task-based model of production (farboodi2021model) characterizing the quality-adjusted output of a firm $i$ at time $t$ as $y_{i,t}=A_{i,t}k_{i,t}^{\alpha}$, where $k_{i,t}^{\alpha}$ represents the units of output of quality $A_{i,t}$ produced by leveraging $k_{i,t}$ units of capital, and $A_{i,t}=g\big((a_{i,t}-a_{i,t}^{opt})^{2}\big)$ is function of the squared distance between $a_{i,t}$ (the production technique of choice for firm $i$ at time $t$) and the optimal technique $a_{i,t}^{opt}=\theta_{t}+\epsilon_{a,i,t}$}, with $\theta_{t}$ and $\epsilon_{a,i,t}$ being the jointly-observed persistent and transitionary components of the optimal technique, and $g$ a strictly decreasing function. This formulation, by means of the introduction of the squared distance between the actual and the optimal production skill, makes the role of technological performance explicit in the determination of production output of automated tasks. Also, this framework complements the treatments illustrated in (autor2025expertise), where a worker is able to execute a task only if the worker’s expertise level is equal or higher than the one required by the task (the hierarchical expertise assumption), and automation deploys below a certain threshold of task complexity (the hierarchical automation assumption), a condition also illustrated in (acemoglu2025simple) as effect of the comparative advantage offered by labor with respect to capital in newer or complex tasks. Interestingly, this task-based model aligns with the idea of generalization difficulty as presented in (chollet2019measure), i.e. the increased complexity introduced by additional levels of uncertainty presenting at execution time with respect to learning time. The compatibility of these seminal models constitutes the foundational ground for our work.
Moving from this referential framework, we conceptualize performance as a stochastic property that emerges from the interaction between task content and skills, conditioned upon local and global generalization difficulty, extending the definition of production technique to include models operated by the human mind. In this context, we focus on cognitive tasks, leaving the opportunity to relax this assumption in further steps. First, we move by abstraction to defining the distance between the actual and optimal skills in terms of loss ($\mathcal{L}$), such that the quality $A_{i,t}$ can be expressed as $A_{i,t}=g\big(\mathcal{L}_{i,t}\big)$, where $\mathcal{L}_{i,t}=f(a_{i,t},a_{i,t}^{opt})$ and $f$ is any appropriate function measuring the absolute gap between the actual and optimal technological skills. By doing this, we open the possibility to consider alternative measures of distance, including divergence or dissimilarity (also, if necessary, multiple measures of loss may be combined). Then, when the range of potential values of $\mathcal{L}_{i,t}$ is known, we can obtain a normalized loss $\mathcal{L}^{\prime}_{i,t}\in[0,1]$ through the normalization process of choice (for example, $\mathcal{L}^{\prime}_{i,t}=[\mathcal{L}_{i,t}-\min(\mathcal{L}_{i,t})]/[\max(\mathcal{L}_{i,t})-\min(\mathcal{L}_{i,t})]$), and a normalized performance $\theta_{i,t}=(1-\mathcal{L}^{\prime}_{i,t})$. This allows to formally transition from the measure of technological distance to the measure of performance, with consequently easier operational deployment. Conceptually, based on this initial framing, our model operates in a simple sequence. First, a skill attempts a task, resulting in a performance ($\theta$) between zero and one. This performance is then translated into a quality-adjusted output ($y$) and then into an economic value ($v$) by applying a skill-specific margin. Critically, poor performance can also trigger a shock, and the associated cost (errerr). The final economic utility ($u$) is simply the resulting value minus any error cost. The following formalisms detail this process.
Given initial conditions at time $t$, our computational pipeline leverages a stochastic process to determine the normalized performance $\theta_{i,c,d,e}$ obtained by the firm $i$ deploying the skill policy $c$ in the execution of a task with difficulty $d$ at time $t+e$, where $e\in[0,E-1]$ is a task-execution epoch. To this end, without sacrificing generality, we implements three skill policies and three levels of task difficulty. The skill policy options are $H$ (human-exclusive, relying on labor allocations only), $M$ (machine-exclusive, i.e. full automation), and $HM$ (combining human and machine skills), while the levels of task difficulty are set to $High$, MedMed and $Low$. The implementation of skill performance is probabilistic, with distributions $P(\theta)_{c,d,e}$ conditioned on $c$, $d$, and $e$. The global generalization difficulty is modeled by different distributions $P(\theta)_{c,d,e}$ for different levels of difficulty $d$, while the effect of local generalization difficulty is inherently defined by each individual probability distribution $P(\theta)_{c,d,e}$.
At each epoch, the performance obtained by the skill policies $H$ or $M$ (denoted as $\theta_{i,c=H,d,e}$ and $\theta_{i,c=M,d,e}$, and more synthetically as $\theta_{H}$ and $\theta_{M}$, respectively) is randomly drawn according to the appropriate probability distribution $P(\theta)_{c,d,e}$. When, instead, the skill policy adopted by the firm is $HM$, the performance computes as $\theta_{HM}=a(\theta_{H},\theta_{M},\gamma_{HM})$, where $a$ is an arbitrary function defining how the individual skills $H$ and $M$ interact, depending also on an augmenting factor $\gamma_{HM}$. Herein, the concept of augmentation evokes amplification (ashby1956introduction), collaboration (4503259) and integration (engelbart1962augmenting), characterizing a human-machine skill policy in ways that go from task decomposition and sub-task delegation to complex collaborative partnerships.
Then, based on the performance $\theta_{i,c,d,e}$, the quality-adjusted production output computes as $y_{i,c,d,e}=g^{\prime}\big(\theta_{i,c,d,e}\big)$, where $g^{\prime}:\theta_{i,c,d,e}\in[0,1]\rightarrow y_{i,c,d,e}\in[0,1]$ is a monotone and non decreasing function (the combined evaluation of $A_{i,t}$ and ${kl}^{\alpha}_{i,t}$ through the function $g^{\prime}$ is a desired behavior of our model, to streamline the formal relationship between task-execution performance and quality-adjusted output). Next, the value of the production output is determined as $v_{i,c,d,e}=max(0,min(1,y_{i,c,d,e}*mc_{c}*\delta_{c,e}))$, where $mc_{c}\in[0,1],\forall c$ is the margin of contribution associated to each skill policy, and $\delta_{c,e}$, which depends from both $c$ and $e$, represents the skill-specific evolution of the margin of contribution over epochs. To this end, $\delta_{c,e}$ is computed by linear interpolation as $\delta_{c,e}=interpolate(1,1+\delta_{c},e,E)$, according to Algorithm 4, where $\delta_{c}=\{\delta_{H},\delta_{HM},\delta_{M}\}$ are configuration parameters determining the increase or decrease of the margin of contribution over time.
The $\delta_{c}$ parameters describe the progress of the firm in walking the learning curve of each skill, as well as the firm’s ability to improve the cost effectiveness of skill development and operations, or the evolution of the market price for the task production under consideration ($\delta_{c}>0$ determines an increase, i.e. a unit of quality-adjusted production increases through epochs, while $\delta_{c}<0$ determines a decrease). Also, we introduce the element of error as a shock conditioned on performance, modeling the economic impact of a low performance not only in terms of reduced output, but also in terms of economic penalty triggered beyond a minimum performance threshold $t_{err}\in[0,1]$. This threshold can be interpreted as a quality control threshold, or a level of safety tolerance. In combination with $t_{err}$, we introduce the parameter $c_{err}\in[0,1]$, which defines the cost of error as a fraction of input in units ($c_{err}$). This represents a variety of phenomena, including the cost of the necessary corrections, the cost of repair for eventual damages, a financial penalty, a reputation damage, or the cost of a product recall. By combining these factors, when $(1-\theta_{i,c,d,e})\geq t_{err}$ we compute $err_{i,c,d,e}=(1-\theta_{i,c,d,e})*c_{err}$, or $err_{i,c,d,e}=0$ otherwise. Finally, the bottom-line economic utility is computed as $u_{i,c,d,e}=v_{i,c,d,e}-err_{i,c,d,e}$.
Figure 1: Schematic representation of the computational flow implemented by the model
The complete process instructed by the model is more formally describe by Algorithm 1, and schematically illustrated in Figure 1. As illustrated in Figure 2, the result of each simulation $s$ is collected through the following sets of values: $\Theta^{s}=\{\theta_{i,c,d,e,k}^{s}\}$, $Y^{s}=\{y_{i,c,d,e,k}^{s}\}$, $V^{s}=\{v_{i,c,d,e,k}^{s}\}$, $ERR^{s}=\{err_{i,c,d,e,k}^{s}\}$, and $U^{s}=\{u_{i,c,d,e,k}^{s}\}$. For notational convenience, we generically refer to these sets as $\Lambda^{s}=\{\lambda_{i,c,d,e,k}^{s}\}$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$ and $\lambda\in\{\theta,y,v,err,u\}$. Additionally, we define the subsets of values characterized by the same skill policy $c_{0}$ as $\Lambda^{s}_{c_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0},\forall i,e,d,k\}\subseteq\Lambda^{s}$, and the subsets of values characterized by the same skill policy / task difficulty pair $(c_{0},d_{0})$ as $\Lambda^{s}_{c_{0},d_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0}\land d=d_{0},\forall i,e,k\}\subseteq\Lambda^{s}_{c_{0}}$. Next, for each set $\Lambda^{s}_{[c],[d]}$ (this compact notation stands for $\Lambda^{s}$, $\Lambda^{s}_{c}$ and $\Lambda^{s}_{c,d}$), we compute the arithmetic mean as $\mu(\Lambda^{s}_{[c],[d]})$, the standard deviation as $\sigma(\Lambda^{s}_{[c],[d]})$, the range as $\rho(\Lambda^{s}_{[c],[d]})=\max(\Lambda^{s}_{[c],[d]})-\min(\Lambda^{s}_{[c],[d]})$, the interquartile range (IQR) as $IQR(\Lambda^{s}_{[c],[d]})$, and the skewness as $SK(\Lambda^{s}_{[c],[d]})$. The descriptive statistics $\mu$, $\sigma$, $\rho$, $IQR$ and $SK$ are also generically denoted as $\omega$, crystallizing $\omega(\Lambda^{s}_{[c],[d]})$ as a general notational handle for the characterization and analysis of results.
Algorithm 1 Model
[…]
10:$K,N,E$
[…]
15:for $k=1$ to $K$ do
16:  for $e=0$ to E1E-1 do
17:    if $e=0$ then
18:     $c\leftarrow randomchoice(C,N,P_{c})$
19:    end if
20:    for $i=0$ to N1N-1 do
21:     $d\leftarrow randomchoice(D,1,P_{d})$
22:     $\alpha_{c,d,e}\leftarrow interpolate(\alpha_{c,d,e=0},\alpha_{c,d,e=E-1},e,E)$
23:     $\beta_{c,d,e}\leftarrow interpolate(\beta_{c,d,e=0},\beta_{c,d,e=E-1},e,E)$
24:     $\theta_{i,c=H,d,e,k}\leftarrow randomsample\big(\alpha_{c=H,d,e},\beta_{c=H,d,e},1)$
25:     $\theta_{i,c=M,d,e,k}\leftarrow randomsample\big(\alpha_{c=M,d,e},\beta_{c=M,d,e},1)$
26:     if $c=H$ then
27:       $\theta_{i,c,d,e,k}\leftarrow\theta_{i,c=H,d,e,k}$
28:     else if $c=M$ then
29:       $\theta_{i,c,d,e,k}\leftarrow\theta_{i,c=M,d,e,k}$
30:     else if $c=HM$ then
31:       $\theta_{i,c,d,e,k}\leftarrow a(\theta_{i,c=H,d,e,k},\theta_{i,c=M,d,e,k},\gamma_{HM})$
32:     end if
33:     $y_{i,c,d,e,k}\leftarrow g^{\prime}(\theta_{i,c,d,e,k})$
34:     $\forall c:\delta_{c,e}\leftarrow interpolate(1,1+\delta_{c},e,E)$
35:     $v_{i,c,d,e,k}\leftarrow y_{i,c,d,e,k}*mc_{c}*\delta_{c,e}$
36:     if $(1-\theta_{i,c,d,e,k})\geq t_{err}$ then
37:       $err_{i,c,d,e,k}\leftarrow(1-\theta_{i,c,d,e,k})*c_{err}$
38:     else
39:       $err_{i,c,d,e,k}\leftarrow 0$
40:     end if
41:     $u_{i,c,d,e,k}\leftarrow v_{i,c,d,e,k}-err_{i,c,d,e,k}$
42:    end for
43:  end for
44:end for
Returns: $\Lambda^{s}=\{\lambda_{i,c,d,e,k}^{s}\}$ with $\lambda\in\{\theta,y,v,err,u\}$, i.e. $\Theta^{s}=\{\theta_{i,c,d,e,k}^{s}\}$, $Y^{s}=\{y_{i,c,d,e,k}^{s}\}$, $V^{s}=\{v_{i,c,d,e,k}^{s}\}$, $ERR^{s}=\{err_{i,c,d,e,k}^{s}\}$, $U^{s}=\{u_{i,c,d,e,k}^{s}\}$ with $s$ denoting the identifying index of the simulation of interest.
Figure 2: Schematic representation of an experiment combining multiple simulations with different input parameters, and the sub-setting/processing of the resulting data
3 Design of experiment
The setup for our Monte Carlo experiments is defined by a combination of experiment-agnostic and experiment-specific configurations. The experiment-agnostic configuration is made of three components: a general setup, an numeric setup, and a categorical setup. The general setup, common to all experiments, specifies $N$, $E$, $K$, $P_{C}$ and $P_{D}$. In our study, we focus on a single firm ($N=1$) walking through ten epochs ($E=10$). With regards to the number of replications, which remarkably impacts the computational effort required by simulations, we set $K=1000$, conditioning the eventual increase of $K$ to the analysis of the margin of error at 95% confidence level at the end of the first round of experiments. At each replication, the firm is randomly assigned a skill policy according to $P_{c}\sim U\{P_{H},P_{HM},P_{M}\}$, and at each epoch the task difficulty is randomly drawn from $P_{d}\sim U\{P_{Low},P_{Med},P_{High}\}$; the choice of uniformly-distributed probabilities is due to realize an equal split of the simulation resources over all the potential cases under analysis.
The numeric setup (two-decimals precision) is obtained by algorithmic selection of the values of $\gamma_{HM}$, $mc_{H}$, $mc_{M}$, $mc_{HM}$, $\delta_{HM}$, $t_{err}$ and $c_{err}$ optimizing the coverage of the parameter space. With regards to the applicable value-ranges for each numeric parameter, we implement plausible calibrations. For $\gamma_{HM}$, we set the range $[1,2]$, limiting the augmenting factor to increase only, up to +100%. For the unitary margins of contribution associated to each skill policy, we set $mc_{H}\in[0.4,0.6]$ and $mc_{M}\in[0.7,0.9]$ to represent a substantial cost advantage for automation. Then, we model the human-machine approach in terms of a dual-skill policy that incurs in a dual-cost structure defined as $mc_{HM}\in[0.1,0.5]$, with the maximum and minimum limits of the range computed as $mc_{HM}^{max}=1-(1-mc_{H}^{max})-(1-mc_{M}^{max})$ and $mc_{HM}^{min}=1-(1-mc_{H}^{min})-(1-mc_{M}^{min})$. For $\delta_{HM}$ we set the range $[0,1]$, thus imposing a limit of +100% to the growth of the margin of contribution of skill policy $HM$ over epochs. Finally, for both $t_{err}$ and $c_{err}$ we set a $[0,1]$ range, to explore the whole space of possibilities. Based on these constrains, the algorithmic selection is operated by combining the results obtained through the Latin Hypercube Sampling method (LHS) (mckay2000comparison) and the Maximin method (johnson1990minimax) (no duplicates found, for a total of 480 unique numeric setups). Before consolidating the algorithmically-optimized values, we investigate the resulting configuration space for $mc_{c}$. To this end, we compute $\Delta mc_{HM}=mc_{HM}-[1-(1-mc_{H})-(1-mc_{M})]$ to represent the gap between the values of $(mc_{H},mc_{HM},mc_{M})$ selected by LHS and Maximin and the ideal assumption that $mc_{HM}$ should account for the cost of both skills. The distribution of $\Delta mc_{HM}$, significantly deviating from normality, shows a 0.00 mean and a 0.14 standard deviation (see Figure 3(a)). By grouping the values of $mc_{H}$ and $mc_{M}$ in ranges of 0.05 granularity and visualizing the means of $\Delta mc_{HM}$ and $mc_{HM}$ by group (see Figures 3(b) and3(c), respectively) we observe that the actual $mc_{HM}$ deployed in simulations has a comparative advantage versus our ideal hypotheses when both $mc_{H}$ and $mc_{M}$ are in the lower portion of their ranges, and a comparative disadvantage when both $mc_{H}$ and $mc_{M}$ are in the upper portion of their ranges. Nevertheless, $mc_{HM}$ is always lower than $mc_{M}$ by construction, and lower than $mc_{H}$ in 93% of the cases resulting from the selection operated by LHS and Maximin . An overview of the values resulting from LHS and Maximin is synthetically illustrated by means of a pairwise representation in Figure 4. To complete the numeric setup, we impose $\delta_{H}=0$ and $\delta_{M}=0$ as a default, representing the case of human and machine skills being already on their plateau of efficiency.
For the categorical setup, we set the options available for the functions $a$ and $g^{\prime}$. With regards to $a$, we implement three options:
$\theta_{HM}=\begin{cases}(\theta_{H},\theta_{M})/2&\text{if }a=mean,\\ \min(1,mean(\theta_{H},\theta_{M})*\gamma_{HM}))&\text{if }a=collaborate,\\ \min(1,max(\theta_{H},\theta_{M})*\gamma_{HM}))&\text{if }a=superpower,\end{cases}$ (1)
where $\theta_{HM}$, $\theta_{H}$ and $\theta_{M}$ are notational simplifications for $\theta_{i,c=HM,d,e}$, $\theta_{i,c=H,d,e}$ and $\theta_{i,c=M,d,e}$, $\forall(i,d,e)$. Herein, meanmean represents the case of simply averaging the abilities of human and machine skills with no augmenting effects, while collaboratecollaborate and superpowersuperpower represent moderate and strong augmentation, respectively (for notational convenience, these options may also be noted as $HM$, $HM+$ and $HM++$). For the function $g^{\prime}$, we deploy two options:
$y=\begin{cases}1/(1+e^{-k(2\theta-1)})&\text{if }g^{\prime}=logistic,\\ [k-\ln(1-\theta)-\ln(\theta)]/2k&\text{if }g^{\prime}=inverse\_logistic,\end{cases}$ (2)
with $k=5$, where $y$ is a notational simplification for $y_{i,c,d,e}$ and $\theta$ stands for $\theta_{i,c,d,e}$. These options reflect two alternative dynamics for the performance-to-output map.
Finally, the experiment-specific setup involves the hypotheses regarding the distributions $P(\theta)_{c,d,e}$. By means of an overall design choice, we model $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$ to benefit from the range constrained to $[0,1]$ by construction (a natural fit for normalized performance $\theta$) and from its flexibility, allowing us to easily model a variety of skill profiles. With regards to the choice of $\alpha_{c,d,e}$ and $\beta_{c,d,e}$, in the absence of micro-data on skill performance we calibrate to align with stylized conditions. In the first experiment (Exp. 1), we model skill performance such that machines have a comparative advantage in low-complexity tasks, while humans have a comparative advantage in high-complexity task, configuring the human skills to improve across all task difficulties over epochs, while machine skills remain static (training at time $t$ only, with no infra-epoch training) -see Figure 5(a). In a second experiment (Exp. 2), we model a higher performance profile for machine skills with respect to Exp. 1 over medium-to-high-difficulty tasks, as illustrated in Figure 5(b). Finally, in a third experimental session (Exp. 3), we replicate the setup of Exp. 2 while imposing a static human performance over epochs. A comprehensive view of the setup parameters $(\alpha,\beta)_{c,d,e=0}$ and $(\alpha,\beta)_{c,d,e=E-1}$ for all the experiments is detailed in Table 1. In short, the setup of Experiment 1 is calibrated to mimic a state of technological advancement in line with the SOTA at the time this manuscript is written, combined with the idea that human skills may have an evolutionary infra-epoch advantage with respect to machine models. Then, with the Experiment 2 we enhance the technology, making artificial skill outperforming human skills whatever the generalization difficulty involved. Finally, with Experiment 3, we penalize human skills by disabling their ability to evolve over epochs. All the experiments are executed on computing instances deploying Python 3.11, generating a total of 86.4 million in-silico task executions.
(a) Distribution of $\Delta mc_{HM}$
(b) Heatmap of the mean of $\Delta mc_{HM}$
(c) Heatmap of the mean of $mc_{HM}$
Figure 3: Analysis of the actual values determined by LHS and Maximin for $mc_{HM}\in[0.1,0.5]$. In 3(c), each heatmap quadrant reports the ideal range for $mc_{HM}$ in the form [min,max] and the mean of the actual values $mc_{HM}$ determined by LHS and Maximin.
(a) LHS method
(b) Maximin method
Figure 4: Pair plot of numeric variables selected for experimental design instructed by optimization methods
(a) Exp. 1
(b) Exp. 2
Figure 5: Experimental setup for the distributions of probabilities $P(\theta)_{c,d,e}$ for all skill policies and difficulty levels at epoch $e=0$ and e=E1e=E-1 (Exp. 1 and 2), representing human skills to evolve through epochs, while machine skills are static. A different mix of the same configurations constitutes the basis for Exp. 3.
Table 1: Parameters $(\alpha,\beta)_{c,d,e}$ for the distributions $P(\theta)_{c,d,e}$
Exp. 1 Exp. 2 Exp. 3
$e=0$ e=E1e=E-1 $e=0$ e=E1e=E-1 $e=0$ e=E1e=E-1
$\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$
$P(\theta)_{c=M,d,e}$ High 2 3 3 2.5 2 3 3 2.5 2 3 2 3
Med 5 5 5 3 5 5 5 3 5 5 5 5
Low 3 2 6 2 3 2 6 2 3 2 3 2
$P(\theta)_{c=H,d,e}$ High 2 4 2 4 6 4 6 4 6 4 6 4
Med 4 2 4 2 8 2.5 8 2.5 8 2.5 8 2.5
Low 10 2 10 2 10 2 10 2 10 2 10 2
4 Results
Contrary to the conventional wisdom that a human-in-the-loop policy is a safe risk-mitigating strategy, our simulations suggest it prone to contextual economic volatility. To quantitatively characterize this effect, we complement the presentation of the plain results of our individual experiments, denoted as Exp. 1, 2, and 3 (see Figure 6(a)), with close-up perspectives focused on situations characterized by high cost of errors ($t_{err}\leq 0.1$ and $c_{err}\geq 0.8$), denoted as Exp. 1*, 2* and 3* (see Figure 6(b)) and by the low or high ability of the organization to walk the experience curve with the human-machine skill policy over epochs ($\delta_{HM}\leq 0.2$ or $\delta_{HM}\geq 0.8$, respectively), denoted as Exp. 1# and 1£ (see Figure 6(c)).
As shown in Figure 6(a), when confronting high-difficulty tasks in the plausible setting of Experiment 1, the human-machine skill policy achieving augmenting effects reveals higher bottom-line utility than both the human-exclusive and the machine-exclusive strategies. However, when failing to produce augmenting effects, the human-machine strategy results the least economically effective option, derailing into negative utility and potentially destroying value instead of building it. This is far from the result one may expect from a reassuring risk-mitigating approach. Interestingly, the synergic (augmenting) human-machine skill policy is the preferable option in high-difficulty tasks across all the experiments when i) the cost of errors is high (see Figure 6(b)) and ii) regardless of the increase rate over epochs of the margin of contribution associated to the human-machine policy (see Figure 6(c)). This suggests that, in situations where poor task execution triggers expensive shocks, combining human and machine skills can be pivotal even when machine skills become extremely powerful. Also, the increase of the margin of contribution associated to the human-machine skill policy over epochs mitigates the erosion of value, while not safely resolving it (see Figure 6(c), with the human-machine skill policy failing to achieve augmentation closing just at break-even). This highlights the importance of augmentation towards the bottom-line economic effectiveness of the human-machine skill policy, even when machine skills do a great job in high-difficulty tasks.
Next, while denoting the emergent critical finding of the all-or-nothing nature of the human-machine skill policy, our simulation data replicate two relevant and established stylized facts from the existing literature, as presented in Section 1. First, all the experimental results illustrated in Figure 6 (central tendencies) align with the notion that automation tends to be the most economically-effective strategy when task difficulty is low-to-medium, underscoring a superiority in repetitive or compartmentalized tasks. On the other hand, the human-exclusive skill policy can be competitive with respect to automation as long as the human skills can better handling generalization difficulty in situation characterized by high cost of errors. These results are not a straightforward consequence of our design assumptions, because hypotheses are formulated on performance and costs/margins, while our considerations pertain to utility as determined through simulation over the chosen parameter space. Equally interesting is that, while performance is a variable that can be controlled by the decision-maker through careful skill selection or programming, the cost of errors is an exogenous factor.
With regards to dispersion, the mean of the distributions $\sigma(U^{s}_{c,d})$ and $\rho(U^{s}_{c,d})$ presented in Figures 7 and 8 highlights a consistently-lower variability for the values in $U^{s}_{c,d}$ over the simulation space for the human-machine skill policy versus both the human-exclusive or machine-exclusive strategies. In contrast, in Figures 9 and 10 we observe that the standard deviation and the range of the distribution $\mu(U^{s}_{c,d})$ for the human-machine skill policy tends to be often slightly higher than than the one realized by the human or machine skills alone. In both the observations, augmenting effects seem not to have a visible impact. Then, our interpretation is that the hybrid skill strategy contributes more to reducing the variability of utility over the spectrum of scenarios under consideration, while slightly but less evidently increasing the variability of $\mu(U^{s}_{c,d})$. This may explain why the combination of human and machine skills can be perceived as a risk-mitigating strategy, because the reduction of variability within a specific context naturally suggests lower risks; however, the variability alone is not sufficient to avoid the risk of being near or below utility break even when the variables describing the context are subject to uncertainty.
A more fine-grained detail of our results is enclosed in the Appendix of this manuscript. In Table 4, 5 and 6 we present the description of the distributions $\omega(\Lambda^{s}_{c,d})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ through their arithmetic mean, standard deviation, range (max-min), and median value. In Figures 14 and 15 we provide an intuitive visualization of the distributions $\mu(U^{s}_{c,d})$ and $\sigma(U^{s}_{c,d})$ by means of overlapping violin and box plots, with breakdown by skill/interaction policy and task difficulty for each experiment and focused perspective.
Figure 6: Mean of the distributions $\mu(U^{s}_{c,d})$ by experiment and by skill/interaction policy
Figure 7: Mean of the distributions $\sigma(U^{s}_{c,d})$ by experiment and by skill/interaction policy
Figure 8: Mean of the distributions $\rho(U^{s}_{c,d})$ by experiment and by skill/interaction policy
Figure 9: Standard deviation of the distributions $\mu(U^{s}_{c,d})$ by experiment and by skill/interaction policy
Figure 10: Range of the distributions $\mu(U^{s}_{c,d})$ by experiment and by skill/interaction policy
5 Discussion
To allow the reader a more precise understanding of the margins of error of $\mu(\Theta^{s}_{[c],[d]})$ and $\mu(U^{s}_{[c],[d]})$ obtained with $K=1000$ as estimates of the true mean ($K\rightarrow\infty$), for each simulation $s$ of all the experiments we compute the estimated MOE at 95% confidence level for $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$ and $\mu(\Lambda^{s}_{c,d})$ with $\Lambda\in\{\Theta,U\}$. Formally, $\forall s$, we compute $\text{MOE}_{95}\big(\mu(\Lambda^{s})\big)=1.96*\sigma(\Lambda^{s})/\sqrt{K^{s}}$, $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c})\big)=1.96*\sigma(\Lambda^{s}_{c})\sqrt{K^{s}_{c}}$, and $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c,d})\big)=1.96*\sigma(\Lambda^{s}_{c,d})\sqrt{K^{s}_{c,d}}$ with $K^{s}=K*E=10000$, and $K^{s}_{c}$ and $K^{s}_{c,d}$ globally and prudentially set to $3000<K*E/|C|\sim 3333$ and $1000<K*E/(|C|*|D|)\sim 1111$, respectively. We acknowledge that, within each simulation $s$, the sets $\Lambda^{s}$, $\Lambda^{s}_{c}$ and $\Lambda^{s}_{c,d}$ pool values that are related to different epochs $e$, thus potentially in relationship with different Beta distributions for skill policies $H$ and $HM$ (this is due to the deterministic interpolation of the parameters $\alpha$ and $\beta$ over epochs for the skill $H$), meaning that the data is heterogeneous and consists of independent samples of smaller size. However, the margins of error computed on the pooled standard deviation represent a conservative upper bound for the granular error that should be selectively computed for each epoch, because the pooled standard deviation combines the actual sampling uncertainty with the additional variability introduced by the parameter interpolation. The distributions of $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{[c],[d]})\big)$ with $\Lambda\in\{\Theta,U\}$ are described in Figure 11 and in Table 2. with mean consistently below 0.01.
(a) $\Lambda^{s}$
(b) $\Lambda^{s}_{c}$
(c) $\Lambda^{s}_{c,d}$
Figure 11: Distributions of the margins of error $\text{MOE}_{95}(\Lambda^{s}_{[c],[d]})$ with $\Lambda\in\{\Theta,U\}$
Table 2: Descriptive statistics for $\text{MOE}_{95}(\Lambda^{s}_{[c],[d]})$ with $\Lambda\in\{\Theta,U\}$
$\text{MOE}_{95}\big(\mu(\Theta^{s})\big)$ $\text{MOE}_{95}\big(\mu(U^{s})\big)$ $\text{MOE}_{95}\big(\mu(\Theta^{s}_{c})\big)$ $\text{MOE}_{95}\big(\mu(U^{s}_{c})\big)$ $\text{MOE}_{95}\big(\mu(\Theta^{s}_{c,d})\big)$ $\text{MOE}_{95}\big(\mu(U^{s}_{c,d})\big)$
mean 0.0044 0.0051 0.0065 0.0067 0.0087 0.0091
std 0.0005 0.0017 0.0016 0.0039 0.0029 0.0059
IQR 0.0009 0.0024 0.0016 0.0053 0.0045 0.0082
max 0.0053 0.0105 0.0096 0.0236 0.0141 0.0368
Q3 0.0048 0.0062 0.0073 0.0089 0.0110 0.0126
Q2 0.0044 0.0050 0.0068 0.0061 0.0091 0.0078
Q1 0.0039 0.0038 0.0057 0.0036 0.0065 0.0044
min 0.0035 0.0018 0.0012 0.0002 0.0000 0.0000
While the structural combined importance of the skill policy $c$ with the augmenting function $a$ in the context of a task presenting generalization difficulty $d$ emerges from the direct analysis of simulation data, it is interesting to better understand the relative importance of the other parameters that take part in the creation of economic value. To this end, we leverage introspection analysis on meta-models learned on our synthetic data by means of i) Permutation Importance analysis and ii) SHapley Additive exPlanations method (SHAP). Permutation Importance analysis (breiman2001random) is a computational method measuring the importance of a feature by determining the increase of the prediction error after the values of that feature are randomly permuted (a larger increase in error denotes a more important feature). In the context of our work, we combine Permutation Importance analysis with cross validation, finally computing mean and standard deviation of the results; also, we aggregate categorical values encoded in one-hot variables in a unique feature, improving interpretability. In contrast, the SHAP method (vstrumbelj2014explaining) (datta2016algorithmic) (lundberg2017unified), named by the acronym of SHapley Additive exPlanations, computes the marginal contribution provided by each feature to the output for any given example, providing local interpretability. Then, global model explanations are obtained by aggregating Shapley values across all examples in the data set, suggesting which features are most influential overall, and how they impact the model’s predictions. As the features $c$ and $a$ have a strong correlation by construction, to avoid compromising the response of permutation analysis and SHAP (strobl2008conditional) (lundberg2017unified), we combine these features in a unique new variable called c_ac\_a, which jointly represents skill policy and augmenting effects.
According to these provisions, we learn the model $\hat{f}:(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})\rightarrow\mu(U_{c,d})$. The absence of correlation between variable is verified by computing the adjusted Generalized Variance Inflation Factor (Adj.GVIF) (fox1992generalized), which results consistently in the range $[1.00,1.01]$ (no correlation). With regards to the statistical learning strategy, we implement a Random Forest regressor (RFr) and a Gradient Boosting regressor (GBr), assessing performance through 10-Fold Cross Validation (metrics: $MSE$ and $R^{2}$, see Table 3). The purpose of measuring the predictive performance is only to make sure that the model is effectively compressing data and learning the information encoded in it, an essential pre-requisite for the interpretation of variable importance.
The results illustrated in Figure 12 and in Figure 13 confirm the remarkable importance of the skill/interaction policy in shaping bottom-line utility, followed by task difficulty. An interesting relative importance is visible for the performance threshold triggering the error shocks ($t_{err}$), the cost of errors ($c_{err}$), the baseline margin of contribution associated to the human-machine skill policy ($mc_{HM}$) and the performance-to-output function $g^{\prime}$, indicating that once the fundamental collaborative dynamic is set, the economic success of the policy is nevertheless influenced by its cost and risk architecture. In contrast, and somehow counter-intuitively, the scalar augmenting factor $\gamma_{HM}$ shows to play a minor role towards utility, suggesting that design of the collaborative process itself (the $a$ function, i.e. the kind of interaction between skills) may be more important than a straightforward performance multiplier. In this sense, a collaborative partnership is something different, not simply something more. When narrowing the meta-learning process to high-difficulty tasks only, the results illustrated in Figures 13(c) and in Figure 13(d) visualize the paradox. The skill/interaction policy c_ac\_a still has the largest impact on utility, but its effect is split: higher-valued red dots (representing the human-machine skill policy with the strongest augmenting effects) drive the highest positive impact on utility, while light-blue dots (representing the non-augmenting human-machine interaction) have a negative impact, confirming that the human-machine skill policy is not a risk-mitigating strategy when the task requires the ability to generalize.
Table 3: Results of 10-Fold Cross Validation over the full set of synthetic data (Exp. 1 + Exp. 2 + Exp. 3)
RFr GBr
Meta-model Avg MSE Avg $R^{2}$ Avg MSE Avg $R^{2}$
$\mu(U_{c,d})=\hat{f}(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$ 0.0071 0.8719 0.0079 0.8587
(a) RFr
(b) GBr
Figure 12: Aggregated Permutation Importance analysis (plots illustrate the mean decrease of $R^{2}$)
(a) RFr
(b) GBr
(c) RFr, $d=High$
(d) GBr, $d=High$
Figure 13: SHAP summary (beeswarm), plots illustrating the mean decrease of $R^{2}$ for the model. Variables are ranked by global importance. Each point represents the SHAP value for a specific instance, with the color indicating the feature’s value. Categorical values (extremes): for task difficulty $d$, blue dots correspond to Low and red dots to High; for skill/interaction policy c_ac\_a, blue dots correspond to $H$ and red dots to $HM++$, with intermediate color scale passing through HMHM (light blue), $M$ (grey), and $HM+$ (light red); for the function $g^{\prime}$, the blue dots represent logistic, while red dots represent inverse_logistic.
6 Limitations
Combining Monte Carlo simulations with machine learning meta-models can provide valuable insights, but limitations must be acknowledged. The primary limitation of this study is that its conclusions are derived from synthetically generated data. The simulation creates an idealized in silico environment whose potential gaps with the real-world should always be considered with the highest regard, as unmodeled interactions could impact the outcomes (law2007simulation). As a consequence, the validity of our findings is contingent upon the fidelity of our model, which is governed by two sets of assumptions: design hypotheses, related to the theoretical mechanics and interactions encoded in the model, and setup hypotheses, concerning the specific characterization of the scenarios of interest through the chosen parameters and probability distributions.
Design hypotheses include, by instance, the specific choice of $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$, and the logics of computation of the economic measures, including the treatments related to error shocks and their cost. Another arbitrary design choice pertains the specific integration of stochastic effects, which in our work are limited to the task-execution performance, leaving the other measures of economic impact to deterministic computation. This can potentially inject a re-scaling effect down through the funnel of economic value, thus making the dynamics of bottom-line measures (for example, utility) more a direct reflection of performance rather than an emergent behavior. With regards to setup hypotheses, while our experimental setup parameters are designed for real-world plausibility, the model’s external validity ultimately requires empirical testing. As a more general principle, results and insights should be intended specific to our explored parameter space. With regards to introspection, we underline that Permutation Importance and SHAP are not meant to reveal a causal relationship between features and outcome, but importance with respect to the predictions computed by the model under consideration. Moreover, we recognize that our definition of utility is focused on economics, thus representing a potentially incomplete picture when seen from another perspective (for example, from the point of view of social sciences).
A final consideration pertains the choice of the number of replications $K$ for each simulation, which trades off the processing effort required by the experiment for the estimated margin of error (MOE) related to $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$, and $\mu(\Lambda^{s}_{c,d})$ computed over the replication space. While a number of replications between 5000 and 10000 should be recommended (mundform2011number), one of our goals is to minimize the computational effort to streamline the time and cost of simulations given an acceptable margin of error. Therefore, the considerations based on the analysis of $\mu(\Lambda^{s}_{[c],[d]})$ are conditioned upon the acceptance of the levels of MOE estimated in Section 5, Figure 11 and Table 2. For the purpose of our study, we acknowledge the current margin of error as sufficient. Should a lower margin of error be required by the reader, the number of replications $K$ would need to be increased. To this end, our code is available with MIT license upon request.
7 Conclusions
The common wisdom in the deployment of artificial skills is built on a comforting assumption: combining humans and machines is a defensive safeguard against worst-case scenarios. This policy is believed to mitigate the risks of full automation while capturing some of its benefits by means of a moderate and cautious compromise. This paper formally challenges this assumption, arguing that it may be not only incorrect but also economically dangerous, and therefore paradoxical.
Our findings reveal that in high-difficulty tasks, the human-machine skill policy becomes a high-risk, high-reward strategy. This is the primary contribution of our work. Jointly, we show that in tasks characterized by high generalization difficulty, when genuine augmentation is not achieved, the human-machine approach frequently becomes the worst of all possible options, and an economic pitfall that actively destroys value by incurring the costs of both human and machine components without a synergistic benefit. However, the risk is mirrored by an equally high potential reward. When surprise is the norm, if genuine augmentation is achieved, the combination of human and artificial skills can unlock the highest economic utility, outperforming either humans or machines alone.
This finding fundamentally highlights the two-folded nature of the managerial challenge of the AI era. The first strategic decision pertains whether or not to blend human and artificial skills. On this subject, our study suggests that the solution to the problem is highly contextual, depending on the skills available, on their performance, on their cost/margin profile and on the difficulty of the task under consideration. The second organizational challenge pertains how to combine human and machine skills towards augmentation when task difficulty is high, because this is the critical success factor conditioning the bottom-line economic results of skill policy decisions. To this end, the default human-in-the-loop paradigm may be insufficient, requiring a focused design effort to develop a collaborative partnership.
In the context of these two challenges, the in-silico framework we propose is a direct-to-purpose tool for ex-ante validation, allowing leaders to test the economic viability of their strategies before costly implementation. By quantifying the economic utility of the skill policies of interest, our experiments provide the foundation for a structured managerial test-before-invest approach supported by the evaluation of the trade-offs between automation, human expertise, and human-machine collaboration. This framework moves beyond static task allocation and instead proposes a dynamic process for effective skill policy evaluation articulated in four stages:
  1. 1.
    Economic problem decomposition (understand the puzzle). Before a policy can be evaluated, the managerial problem must be formally decomposed. This requires to i) rigorously define tasks and to classify them by generalization difficulty, ii) model task-execution performance in terms of distribution of probability for all available skills relevant to the task, and iii) map the economic utility of outcomes, including the cost of errors;
  2. 2.
    Focus on synergistic augmentation (design 1+1=3). Our findings suggest that, confronting high-difficulty tasks, the human-machine policy is a high-variance strategy whose success is contingent on achieving genuine augmentation. This implies that managers must shift their focus from deciding allocations to designing collaborative workflows that actively engage human and machine skills in a process that can allow a better handling of generalization difficulty;
  3. 3.
    Policy simulation (search at scale). Given the complex, emergent nature of skill-task interactions, a priori analytical solutions are often intractable. The methodology presented in this paper offers means of searching at scale. By calibrating the framework with the parameters and methods determined at stage 1 and 2, decision-makers can conduct in-silico experiments to explore the economic outcome of competing skill policies under various conditions, identifying the option that yields the maximum expected utility, while mitigating downside risks;
  4. 4.
    Dynamic policy refinement (act and adapt). The method we propose is not a static, one-time tool but a mechanism for adaptive management. Data gathered from real-world executions of tasks should be used to continuously re-calibrate the model. This creates a dynamic feedback loop, enabling organizations to refine and adapt their skill policies in response to evolving technological capabilities and changing economic conditions.
Naturally, this study has limitations: while our framework is built with plausibility in mind, it is not a direct field study. Therefore, it should be viewed as a computational thought experiment designed to test the logical consequences of our plausible assumptions, requiring future empirical testing for final validation. Additionally, future work could extend this model to include dynamic learning and the impact of different organizational structures.
This paper serves as a formal economic challenge to the myth of human-machine integration as a safe harbor. It is a call to move beyond wishful thinking and toward a critical, evidence-based approach for deciding when (and when not) to combine human and artificial skills, improving the organization’s competitiveness and the economic effectiveness of AI initiatives through highly-effective skill policy decisions. Finally, we hope this work contributes to seeing a thoughtful combination of human and machine skills not merely as an ethical consideration, but as an economically-justified strategy.
8 Future developments
Empirical validation would radically strengthen this work, and should be the next-step priority. Additionally, modeling the unitary margins of contribution and the cost of error as probability distributions to reflect the variability of cost structures may be considered. Similarly, exploring other setups for the parameters characterizing the probability distributions of performance may be a promising option, together with considering a variety of classes of distributions. Moreover, the evolution of skills through linear interpolation may be refined to consider the nuances of skill learning and adaptation. Also, modeling the evolution of skills as a non-linear function of cumulative experience, or data generated, is an interesting perspective, in line with the provisions illustrated in (farboodi2021model), where data is a defined as a by-product of production potentially enabling and feeding a feedback loop. Finally, future designs may specifically model dynamic training for machine skills.
9 Acknowledgments
The author declares no conflict of interest and no competing interests at the time this article is published.
Appendix A Appendix
Table 4: Statistical characterization of the distributions $\omega(\Lambda^{s}_{c,d})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ by mean, standard deviation, range and median value (Exp. 1, 2 and 3)
Table 5: Statistical characterization of the distributions $\omega(\Lambda^{s}_{c,d})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ by mean, standard deviation, range and median value (Exp. 1*, 2* and 3*)
Table 6: Statistical characterization of the distributions $\omega(\Lambda^{s}_{c,d})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ by mean, standard deviation, range and median value (Exp. 1# and 1£)
(a) $\mu(U^{s}_{c,d})$, Exp. 1
(b) $\mu(U^{s}_{c,d})$, Exp. 2
(c) $\mu(U^{s}_{c,d})$, Exp. 3
(d) $\mu(U^{s}_{c,d})$, Exp. 1*
(e) $\mu(U^{s}_{c,d})$, Exp. 2*
(f) $\mu(U^{s}_{c,d})$, Exp. 3*
(g) $\mu(U^{s}_{c,d})$, Exp. 1#
(h) $\mu(U^{s}_{c,d})$, Exp. 1£
Figure 14: Distributions of the $\mu(U^{s}_{c,d})$, with breakdown by skill policy, task difficulty, and augmenting effect $a$ (overlapping violin and box plots)
(a) $\sigma(U^{s}_{c,d})$, Exp. 1
(b) $\sigma(U^{s}_{c,d})$, Exp. 2
(c) $\sigma(U^{s}_{c,d})$, Exp. 3
(d) $\sigma(U^{s}_{c,d})$, Exp. 1*
(e) $\sigma(U^{s}_{c,d})$, Exp. 2*
(f) $\sigma(U^{s}_{c,d})$, Exp. 3*
(g) $\sigma(U^{s}_{c,d})$, Exp. 1#
(h) $\sigma(U^{s}_{c,d})$, Exp. 1£
Figure 15: Distributions of the $\sigma(U^{s}_{c,d})$, with breakdown by skill policy, task difficulty, and augmenting effect $a$ (overlapping violin and box plots)
Algorithm 2 Function $randomchoice:(X,n,P(X))\rightarrow choice$
[…]
2:$n$
[…]
Return: choicechoice $\triangleright$ A vector of length $n$
Algorithm 3 Function $randomsample:(\alpha,\beta,n)\rightarrow sample$
[…]
3:$n$
[…]
Return: samplesample $\triangleright$ A vector of length $n$
Algorithm 4 Function $interpolate:(param\_start,param\_end,cur\_step,tot\_steps)\rightarrow cur\_param$
1:param_startparam\_start
2:param_endparam\_end
3:cur_stepcur\_step
4:tot_stepstot\_steps
5:if $total\_steps\leq 1$ then
6:  $cur\_param\leftarrow param\_start$
7:else
8:  $cur\_param\leftarrow param\_start+(param\_end-param\_start)*cur\_step/(tot\_steps-1)$
9:end if
Return: cur_paramcur\_param